\begin{thebibliography}{lll}
\bibitem{DW} L. van den Dries, A. J. Wilkie. Gromov's Theorem on groups of polynomial
growth and elementary logic, J. Algebra 89 (1984), 349--374.

\bibitem{D} C. Dru\c tu. Quasi-isometry invariants and asymptotic cones, Int. J. Alg.
Comp. 12 (2002), 99--135.

\bibitem{DS} C. Dru\c tu and  M. Sapir. Tree-graded spaces and asymptotic
cones of groups. With an appendix by D. Osin and M. Sapir.
Topology 44  (2005),  no. 5, 959--1058.


\bibitem{G1} M. Gromov. Groups of polynomial growth and expanding maps, Inst.
Hautes \'Etudes Sci. Publ. Math. No. 53 (1981), 53--73.

\bibitem{G2} M. Gromov. Asymptotic Invariants of Infinite Groups, Geometric Group
Theory (vol. 2), G. A. Niblo, M. A. Roller (eds), Proc. of the Symposium
held in Sussex, LMS Lecture Notes Series 181, Cambridge University Press
1991.


\bibitem{H} E. Hrushovski. Stable group theory and approximate subgroups, J. Amer.
Math. Soc. 25 (2012), no. 1, 189--243.

\bibitem{Pan} P. Pansu,  M\'etriques de Carnot-Carath\'eodory quasiisom\'etries des espaces sym\'etriques de rang un. Ann. of Math. 129, 1--60 (1989).

\bibitem{P} F. Point. Groups of polynomial growth and their associated metric spaces,
J. Algebra 175 (1995), 105--121.

%\bibitem{OOS} A. Ol'shanskii, D. V. Osin, M. Sapir. Lacunary hyperbolic groups, Geom.
%Topol. 13 (2009), 2051--2140.

\bibitem{SBMS} M. Sapir.
Asymptotic invariants, complexity of groups and related problems. \emph{Bull. Math. Sci.} 1 (2011), no. 2, 277--364.

\bibitem{Sis} A. Sisto.
Separable and tree-like asymptotic cones of groups. arXiv:1010.1199 .

\bibitem{TV} S. Thomas, B. Velickovic. Asymptotic cones of finitely generated groups,
Bull. London Math. Soc. 32 (2000), no. 2, 203--208.
\end{thebibliography}

\documentclass[11pt, makeidx]{amsart}

\usepackage{enumerate}
\usepackage{amssymb,amsfonts,amsthm}
\usepackage{srcltx}

\usepackage{float}

\usepackage[OT2,OT1]{fontenc}
\newcommand\cyr{%
\renewcommand\rmdefault{wncyr}%
\renewcommand\sfdefault{wncyss}%
\renewcommand\encodingdefault{OT2}%
\normalfont
\selectfont}
\DeclareTextFontCommand{\textcyr}{\cyr}

\usepackage{color}
\makeindex

\makeatletter
\def\sssub{\@startsection{paragraph}{4}}

\renewcommand\paragraph{\@startsection{paragraph}{4}{\z@}{1.25ex}{0.0001pt}{\normalfont\normalsize\em}}

\numberwithin{paragraph}{subsubsection}

\newcommand{\R}{{\mathbb R}}

\newcommand{\N}{\mathbb{N}}

\newcommand{\Con}{{\mathrm{Con}}}

\newcommand{\dist}{{\mathrm{dist}}}

\newtheorem{theorem}{Theorem}

\theoremstyle{definition}
\newtheorem{df}[theorem]{Definition}

\newcommand{\iv}{^{-1}}

\newenvironment{dedication}
    {\vspace{3ex}\begin{quotation}\begin{center}\begin{em}}
    {\par\end{em}\end{center}\end{quotation}}
\begin{document}

\title[On groups with locally compact asymptotic cones]{On groups with locally compact asymptotic cones} \author{Mark Sapir}\thanks{The research was supported in part by NSF and BSF
grants} \address{Department of Mathematics, Vanderbilt
University, Nashville, TN 37240, U.S.A.} \email{m.sapir@vanderbilt.edu}
\subjclass[2000]{{Primary 20F65; Secondary 20F69, 20F38, 22F50}}
\keywords{asymptotic cone, locally compact, group}

\maketitle
\begin{dedication}
Dedicated to Stuart Margolis' 60s birthday
\end{dedication}
\bigskip

\begin{abstract} We show how a recent result of Hrushovsky \cite{H} implies that if an asymptotic cone of a finitely generated group is locally compact, then the group is virtually nilpotent.
\end{abstract}

\bigskip

Let $G$ be a group generated by a finite set $X$.
Then $G$ can be considered as a metric space where $\dist(g,h)$ is the length of a shortest word on $X\cup X\iv$ representing
$g\iv h$. Let $\omega$ be a non-principal  ultrafilter on $\N$, i.e. a function from the set of subsets $P(\N)$ to $\{0,1\}$ with $\omega(A\cup B)=\omega(A)+\omega(B)$ if $A\cap B=\emptyset$, $\omega(\N)=1$ and $\omega(A)=0$ for every finite set $A$.

Note that if $A\subseteq \N$, then $\omega(A)+\omega(\N\setminus A)=1$ because $\N$ is a disjoint union of $A$ and $\N\setminus A$. For every $A\subseteq B\subseteq \N$, we have  $\omega(A)
\le \omega(B)$ because $B$ is a disjoint union of $B\setminus A$ and $A$. For every two subsets $A,B\subseteq \N$, $\omega(A\cup B)\le \omega(A)+\omega(B)$ because $A\cup B$ is a disjoint union of $A\setminus B\subseteq A$ and $B$. Also if $\omega(A)=\omega(B)=1$, then $\omega(A\cap B)=1$ because $A\cap B=\N\setminus ((\N\setminus A)\cup (\N\setminus B))$.  For every property $P$ of natural numbers, we say that $P$ is true $\omega$-almost surely if the set $S$ of numbers with property  $P$ satisfies $\omega(S)=1$. Thus is $\omega$-almost surely every natural number satisfies $P$ and $\omega$-almost surely every natural number satisfies $Q$, then $\omega$-almost surely every natural number satisfies $P$ and $Q$. We are going to use this property of ultrafilters several times later without reference.

For every sequence of real numbers $r_i\ge 0$ one can define the limit $\lim_\omega r_i$ as the (unique) number $r\in \R\cup\{\infty\}$ such that for every $\epsilon$ we have $|r_i-r|<\epsilon$ $\omega$-almost surely for all $i\in \N$ (see \cite{D}).

Choose a sequence of \emph{scaling constants} $d_n>0$ such that $\lim_\omega d_n =\infty$ and define a pseudo-metric $\overline\dist$ on the Cartesian power $G^\N$ as $$\overline\dist((u_i),(v_i))=\lim_\omega \frac{\dist(u_i,v_i)}{d_i}.$$ Let $G^\N_b$ be the connected component of the sequence $\bar 1=(1,1,\ldots)\in G^\N$. Thus $G^\N_b$ consists of all the elements from $G^\N$ at finite distance from $\bar 1$. Let $\sim$ be the equivalence relation $(u_i)\sim (v_i)$ if and only if $\overline\dist((u_i),(v_i))=0$. Then $\overline\dist$ induces a metric $\dist^\omega$ on $G^\N_b/\sim$, and $G^\N_b/\sim$ with this metric is called the \emph{asymptotic cone} of $G$ corresponding to the ultrafilter $\omega$ and the sequence of scaling constants $(d_n)$, denoted $\Con^\omega(G,(d_i))$ (for more details see \cite{D,DS,SBMS}). Elements of $\Con^\omega(G,(d_n))$ corresponding to sequences $(y_i)\in G^\N$ will be denoted by $(y_i)^\omega$. Note that $G^\N_b$ is a group which acts on $\Con^\omega(G,(d_i))$ transitively by isometries: $(g_i)\cdot (u_i)^\omega=(g_iu_i)^\omega$ (see \cite{D,SBMS}).

Asymptotic cones were first explicitly used by van den Dries and Wilkie \cite{DW} to prove the celebrated theorem of Gromov \cite{G1} that every finitely generated group of polynomial growth is virtually nilpotent (for more applications see \cite{G2,D,DS,SBMS}). One can easily deduce from \cite{DW} that if all asymptotic cones of a finitely generated group are proper (i.e. all closed balls are  compact), then $G$ has polynomial growth and hence is virtually nilpotent.

A finitely generated group can have several non-homeomorphic asymptotic cones \cite{TV}. In fact there are finitely generated groups with continuum non-$\pi_1$-equivalent asymptotic cones \cite{DS}. On the other hand, every virtually nilpotent finitely generated group has only one asymptotic cone up to isometry by Pansu \cite{Pan}.
In \cite{DW}, van den Dries and Wilkie asked whether local compactness of just one asymptotic cone of $G$ implies that $G$ is virtually nilpotent.

F. Point proved \cite{P} that if one asymptotic cone of $G$ is locally compact and has finite Minkovsky dimension, then $G$ is virtually nilpotent.

Here we will show how to use a recent result of Hrushovsky \cite[Theorem 7.1]{H} to answer the question of van den Dries and Wilkie affirmatively. To formulate Hrushovsky's result we need the notion of an approximable subgroup.

\begin{df}\label{d:a} Let $k\in \N$. A finite subset
$X$ of a group $G$  is said to be a $k$-approximate subgroup if $1 \in X$, $X = X\iv,$
and $XX$ is contained in the union of $k$ left
cosets $gX$ of $X$. (Hrushovsky uses right cosets but it does not affect the results because $x\mapsto x\iv$ is an anti-isomorphism for every group.)
\end{df}

\begin{theorem}[Hrushovsky \cite{H}] \label{t:71}
Let
$G$ be
a finitely generated group, $k\in N$.
Assume that $G$ has a collection $\{F_i, i\in \N\}$ of
$k$-approximate subgroups such that any finite subset $F \subset G$ is contained in a one of the $F_i$. Then $G$ is virtually nilpotent.
\end{theorem}

\begin{theorem} \label{t:1} If one asymptotic cone of a finitely generated group $G$ contains a compact closed ball of positive radius, then $G$ is virtually nilpotent.
\end{theorem}

\proof Suppose that a closed ball of radius $\epsilon>0$ in $C=\Con^\omega(G, (d_i))$ is compact. Since $C$ admits a transitive group of isometries, we can assume that this is the ball  $B_C(\epsilon,\bar 1)$ of radius $\epsilon$ around $\bar 1$. Since $B_C(\epsilon,\bar 1)$ is compact, it is contained in the union of a finite number of balls $B_C(\frac{\epsilon}{4}, x_i)$, $x_i\in B_C(\frac78\epsilon,\bar 1)$, $i=1,\ldots,k$.  Let $x_i=(g_{i,j})^\omega$, $i=1,\ldots,k$, $j\in \N$. Note that by the definition of asymptotic cone, for every $i=1,\ldots,k$, $\dist(g_{i,j},1)< \epsilon d_j$ $\omega$-almost surely (since $\frac78<1$).

For every $j\in \N$ let $2p_j$ be the smallest even natural number which is greater than $\epsilon d_j$.
Then $|2p_j-\epsilon d_j|< 2$ for every $j$, whence $\lim_\omega p_j=\infty.$

For every $i=1,\ldots,k$ and $j\in \N$ consider the balls $B_G(2p_j,1)$ and $B_G(p_j,g_{i,j})=g_{i,j}B_G(p_j,1)$ in $G$ (viewed as a metric space).

We claim that $\omega$-almost surely for every $j$, the ball $B_G(2p_j,1)$ is covered by the balls $B_G(p_j,g_{i,j})$, $i=1,\ldots,k$. Indeed, let $S$ be the set of natural numbers $j$ for which this is true. Suppose that  $\omega(S)=0$.  Then $\omega(\N\setminus S)=1$. For each $j\in \N\setminus S$ let $y_j$ be a point in $B_G(2p_j,1)$ which does not belong to any of the balls $B_G(p_j,g_{i,j})$. Consider any sequence $(z_i)\in G^\N$ such that $z_j=y_j$ for all $j\in \N\setminus S$. Since $\omega(\N\setminus S)=1$, $z=(z_i)^\omega$ is at distance at least $\lim_\omega \frac{p_j}{d_j}=\frac{\epsilon|}2$ from $(g_{i,j})^\omega=x_i$, $i=1,\ldots,k$. Therefore $z$ does not belong to any of the balls $B_C(\frac{\epsilon}{4},x_i)$. Hence $z$ does not belong to $B_C(\epsilon,\bar 1)$. But $\dist^\omega(z,\bar 1)=\lim_\omega \frac{\dist(z_j,1)}{d_j}\le \epsilon$, a contradiction.

Thus $\omega$-almost surely for every $j$, the ball $B_G(2p_j, 1)$ is contained in the union $$\bigcup_{i=1}^k g_{i,j}B_G(p_j,1).$$ But $B_G(2p_j,1)$ is equal to the product $B_G(p_j,1)B_G(p_j,1)$ in $G$. Hence $B_G(p_j, 1)$ is a $k$-approximable subgroup of $G$ $\omega$-almost surely for every $j$. Since $G$ is finitely generated and $\lim_\omega p_j=\infty$, every finite subset of $G$ is contained in one of the balls $B_G(p_j,1)$, so it remains to apply Theorem \ref{t:71}.
\endproof

{\bf Acknowledgement} This note was inspired by the paper \cite{Sis} where an attempt to answer the question of van den Dries and Wilkie was made. I am grateful to Terry Tao who suggested that the answer to that question should follow from \cite{H}.

\end{document}